# Minimum-norm solutions of the non-symmetric semidefinite Procrustes problem

### v1

@@ Abstract
Given two matrices $X,B\in\mathbb{R}^{n\times m}$ and a set $\mathcal{A}\subseteq\mathbb{R}^{n\times n}$, a Procrustes problem consists in finding a matrix $A\in\mathcal{A}$ such that the Frobenius norm of $AX-B$ is minimized. When $\mathcal{A}$ is the set of the matrices whose symmetric part is positive semidefinite, we obtain the so-called non-symmetric positive semidefinite Procrustes (NSPDSP) problem. The NSPDSP problem arises in the estimation of compliance or stiffness matrix in solid and elastic structures. If $X$ has rank $r$, Baghel et al. (Lin. Alg. Appl., 2022) proposed a three-step semi-analytical approach: (1) construct a reduced NSPDSP problem in dimension $r\times r$, (2) solve the reduced problem by means of a fast gradient method with a linear rate of convergence, and (3) post-process the solution of the reduced problem to construct a solution of the larger original NSPDSP problem. In this paper, we revisit this approach of Baghel et al. and identify an unnecessary assumption used by the authors leading to cases where their algorithm cannot attain a minimum and produces solutions with unbounded norm. In fact, revising the post-processing phase of their semi-analytical approach, we show that the infimum of the NSPDSP problem is always attained, and we show how to compute a minimum-norm solution. We also prove that the symmetric part of the computed solution has minimum rank bounded by $r$, and that the skew-symmetric part has rank bounded by $2r$. Several numerical examples show the efficiency of this algorithm, both in terms of computational speed and of finding optimal minimum-norm solutions.
@@ 1 Introduction
A structured mapping problem associated to the set $\mathcal{A}\subseteq\mathbb{R}^{n\times n}$ consists in finding a matrix $A\in\mathcal{A}$ such that $AX=B$, where $X,B\in\mathbb{R}^{n\times m}$ are two fixed matrices; see, e.g., [1, 16] and the references therein. The corresponding Procrustes problem aims to solve a mapping structured problem when it is inconsistent, that is, when there is no matrix $A\in\mathcal{A}$ such that $AX=B$, leading to
$\inf_{A\in\mathcal{A}}\|AX-B\|,$ (1)
[…]
@@
where $\|\cdot\|_{F}$ denotes the Frobenius norm. These problems are both convex, since they minimize a convex function over a convex set. While ($\mathcal{P}$) is well-known in the literature and has been studied in many works [5, 20, 14, 7, 12], ($\mathcal{NP}$) is more recent and has been considered in [14, 13, 4]. Both PSDP and NSPDSP problems arise in the estimate of local compliance matrices during deformable object in various engineering applications; see [13, 4]. The compliance matrix represents the Green function matrix that corresponds to the unknown $A$ in ($\mathcal{NP}$). The estimation of $A$ is based on the observations of the forces, stored in the columns of $X$, and their associated displacements, stored in the columns of $B$, for a total of $m$ measurements. When the deformable object is a passive object, that is, it does not generate energy in deformation, the physical properties of the model impose that $p^{\top}Ap\geq 0$, where $p$ is the point load, which is implied by $A\in\mathcal{N}^{n}_{\succeq}$. Thus, when estimating $A$ under noisy measurements, it is crucial to impose the constraint in the optimization problem.
@@ 1.1 Contributions and outline of the paper
For both ($\mathcal{P}$) and ($\mathcal{NP}$), the original $n\times m$ dimensional problem can be reduced to a problem of dimension $r\times r$, where $r$ is the rank of $X$; see [7] for ($\mathcal{P}$), and [4] for ($\mathcal{NP}$). This is referred to as a semi-analytical approach because the smaller $r\times r$ dimensional problem still needs to be solved with iterative methods, and [7, 4] relied on an optimal first-order method with linear convergence. It is important to note that the solution of ($\mathcal{P}$) is not always attained [7]; see Section 2.1 where we recall this result. The approach in [4] follows closely that of [7], and concludes that the solution of ($\mathcal{NP}$) is also not always attained. However, it turns out this result is not correct, unless one uses an additional constrained within the NSPSDP problem, which is not explicitly stated in [4].
@@
The paper is organized as follows. In Section 2, we recall the semi-analytical methods proposed in [7] for solving ($\mathcal{P}$) and in [4] for solving ($\mathcal{NP}$). We discuss in detail the main issue of the latter approach, namely the introduction of an unnecessary additional constraint in the problem, which leads to non-optimal and unbounded solutions. In Section 3, we show how to fix this issue, and prove that the infimum of the NSPSDP problem is always attained, and we discuss how to compute the minimum-norm solution and show its low-rank properties. In Section 4, we develop an algorithm relying on the theoretical results from Section 3. Section 5 shows the effectiveness of this new algorithm compared to that of [4] in several numerical examples.
@@ 2.1 The PSDP problem
Before doing this, let us recall an important result about PSD characterization that will be used throughout the paper.
@@
[20, Theorem 2.2] Let $X,B\in\mathbb{R}^{n\times m}$ and assume that $X$ has rank $n$. Then the infimum of problem ($\mathcal{P}$) is attained for a unique solution $A_{\star}\in\mathcal{S}_{\succeq}^{n}$.
@@
    Otherwise the infimum is not attained, and we can construct a solution $A_{\star}^{(\varepsilon)}$ with minimal rank, minimum Frobenius norm and minimum spectral norm, such that
@@
    for any $\varepsilon>0$ sufficiently small; see [7] for the explicit formula and Theorem 2.5 for an analogous expression of $\varepsilon$ in the case of the NSPSDP problem.
@@
We refer to [7, Theorem 1] and [7, Corollary 1] for the proof. ∎
@@ 2.2 The NSPSDP problem
The result of Theorem 2.3 for problem ($\mathcal{P}$) has been adapted by Baghel et al. [4] for ($\mathcal{NP}$), but the approach they proposed contains some incorrect statements. Before showing the issues in the main result of [4], we start by recalling an analogous result to Lemma 2.2 for the NSPDSP problem, which guarantees a unique solution when $X$ has full row rank.
@@
[14, Theorem 2.4.6] Let $X,B\in\mathbb{R}^{n\times m}$ and assume that $X$ has rank $n$. Then the infimum of problem ($\mathcal{NP}$) is attained for a unique solution $A_{\star}\in\mathcal{N}_{\succeq}^{n}$.
@@
which is different than the original NSPDSP problem ($\mathcal{NP}$). Although the infimums of ($\mathcal{NPC}$) and ($\mathcal{NP}$) coincide, the optimal solution of ($\mathcal{NPC}$) might not be attained in some cases, see Theorem 2.5 below, while that of ($\mathcal{NP}$) always is; see Theorem 3.1 in the next section.
@@
    Since the infimum of ($\mathcal{NP}$) and ($\mathcal{NPC}$) coincide, see (4), $A_{\star}$ is also an optimal solution of ($\mathcal{NP}$).
@@
    Otherwise the infimum is not attained, and we can construct a solution $A_{\star}^{(\varepsilon)}$ such that
    $\|A_{\star}^{(\varepsilon)}X-B\|^{2}_{F}\leq\inf_{A\in\mathcal{N}_{\succeq}^{n}}\|AX-B\|^{2}_{F}+\varepsilon.$
@@
$A_{\star}=U\left(\left(\begin{matrix}\hat{H}_{11}&\frac{1}{2}Z\\ \frac{1}{2}Z&K\\ \end{matrix}\right)+\left(\begin{matrix}\hat{S}_{11}&-\frac{1}{2}Z\\ \frac{1}{2}Z&R\\ \end{matrix}\right)\right)U^{\top}=U\left(\begin{matrix}\hat{A}_{11}&0\\ Z&K+R\\ \end{matrix}\right)U^{\top},$
@@
Besides solving a different NSPDSP problem with the additional constraint $A\in\mathcal{N}_{\succeq}^{U\mathcal{C}U^{\top}}$, the issue of the method by Baghel et al. concerns the norm of the solution found when the infimum of ($\mathcal{NPC}$) is not attained, that is, when $X$ has low-rank and $\ker(\hat{H}_{11})\neq\{0\}$ is not contained in $\ker(Z)$. In that case, the matrix $K^{(\varepsilon)}\succeq\frac{1}{4}Z(\hat{H}^{(\varepsilon)}_{11})^{-1}Z^{\top}$ used to define $A_{\star}^{(\varepsilon)}$ has a very large norm, going to infinity as $\varepsilon$ goes to zero. This leads to a huge norm solution, as it will be shown in the numerical examples.
3 A new approach for the NSPDSP problem
@@
the infimum of the NSPDSP problem ($\mathcal{NP}$) is always attained.
We characterize the set of minimizers of the problem and we determine the family of the solutions for the original NSPDSP problem.
@@
$\ker(\hat{H}_{11})\subseteq\ker(Z+M),\qquad N\succeq\frac{1}{4}(Z+M)H_{11}^{\dagger}(Z+M)^{\top}.$
@@
Now that we have introduced a new family of solutions for the NSPDSP problem ($\mathcal{NP}$), we provide a simplified characterization of the matrix family $\mathcal{F}$, in which the constraint concerning the kernels in (9) is removed. In this way it will be also easier to highlight low-norm and low-rank properties of matrices in the family.
@@
With the same notation as in Theorem 3.1, the set $\mathcal{F}$ that can be used to characterize optimal solutions of ($\mathcal{NP}$), see (10), can be written as
@@ 3.1 Minimum-norm solution for the NSPSDP problem
$A=U\begin{pmatrix}\hat{A}_{11}&(YW_{1}^{\top}-Z)^{\top}\\ Z&N+G\\ \end{pmatrix}U^{\top}\in U\mathcal{F}U^{\top}.$
Let us introduce the differentiable function
$f(Y)=\|YW_{1}^{\top}-Z\|_{F}^{2}+\frac{1}{16}\left\|Y\Lambda^{-1}Y^{\top}\right\|_{F}^{2},\qquad Y\in\mathbb{R}^{(n-r)\times s}.$ (11)
Note that $f$ depends on $X$ and $B$, since $W_{1}$, $Z$ and $\Lambda$ do. The following result provides the minimum Frobenius norm solution of problem ($\mathcal{NP}$) and it also shows its low-rank properties.
@@
With the same notation as Theorem 3.1 and Lemma 3.2, the minimum Frobenius norm solution of ($\mathcal{NP}$) is
$A_{\star}=U\begin{pmatrix}\hat{A}_{11}&(Y_{\star}W_{1}^{\top}-Z)^{\top}\\ Z&\frac{1}{4}Y_{\star}\Lambda^{-1}Y_{\star}^{\top}\\ \end{pmatrix}U^{\top},$
where
$Y_{\star}=\operatorname*{arg\,min}_{Y\in\mathbb{R}^{(n-r)\times s}}\|YW_{1}^{\top}-Z\|_{F}^{2}+\frac{1}{16}\|Y\Lambda^{-1}Y^{\top}\|_{F}^{2}=\operatorname*{arg\,min}_{Y\in\mathbb{R}^{(n-r)\times s}}f(Y).$
@@
Since the Frobenius norm is unitary invariant and using the characterization of Lemma 3.2, we obtain
@@
The function $f(Y)$ defined in (11) is continuous with bounded level sets, since for any $\rho>0$
$f(Y)\leq\rho\quad\Rightarrow\quad\|Y\|_{F}=\|YW_{1}^{\top}\|_{F}\leq\|YW_{1}^{\top}-Z\|_{F}+\|Z\|_{F}\leq\sqrt{\rho}+\|Z\|_{F},$
and hence the infimum in the right hand side of (12) is attained. Since $f(Y)$ is the sum of a strictly convex function (since $W_{1}$ has orthogonal columns) and a convex function, its minimum is unique and so the first statement follows. The symmetric part of $A_{\star}$ is

### v2

@@ Abstract
Given two matrices $X,B\in\mathbb{R}^{n\times m}$ and a set $\mathcal{A}\subseteq\mathbb{R}^{n\times n}$, a Procrustes problem consists in finding a matrix $A\in\mathcal{A}$ such that the Frobenius norm of $AX-B$ is minimized. When $\mathcal{A}$ is the set of the matrices whose symmetric part is positive semidefinite, we obtain the so-called non-symmetric positive semidefinite Procrustes (NSPSDP) problem. The NSPSDP problem arises in the estimation of compliance or stiffness matrix in solid and elastic structures. If $X$ has rank $r$, Baghel et al. (Lin. Alg. Appl., 2022) proposed a three-step semi-analytical approach: (1) construct a reduced NSPSDP problem in dimension $r\times r$, (2) solve the reduced problem by means of a fast gradient method with a linear rate of convergence, and (3) post-process the solution of the reduced problem to construct a solution of the larger original NSPSDP problem. In this paper, we revisit this approach of Baghel et al. and identify an unnecessary assumption used by the authors leading to cases where their algorithm cannot attain a minimum and produces solutions with unbounded norm. In fact, revising the post-processing phase of their semi-analytical approach, we show that the infimum of the NSPSDP problem is always attained, and we show how to compute a minimum-norm solution. We also prove that the symmetric part of the computed solution has minimum rank bounded by $r$, and that the skew-symmetric part has rank bounded by $2r$. Several numerical examples show the efficiency of this algorithm, both in terms of computational speed and of finding optimal minimum-norm solutions.
Keywords. Procrustes problem, positive semidefinite matrices, minimum-norm solutions, minimum-rank solutions.
AMS subject classification. 15A29, 65F20, 90C22.
@@ 1 Introduction
A structured mapping problem associated to the set $\mathcal{A}\subseteq\mathbb{R}^{n\times n}$ consists in finding a matrix $A\in\mathcal{A}$ such that $AX=B$, where $X,B\in\mathbb{R}^{n\times m}$ are two fixed matrices; see, e.g., [1, 18] and the references therein. The corresponding Procrustes problem aims to solve a mapping structured problem when it is inconsistent, that is, when there is no matrix $A\in\mathcal{A}$ such that $AX=B$, leading to the following optimization problem
$\inf_{A\in\mathcal{A}}\|AX-B\|,\quad\text{ where $\|\cdot\|$ is some norm. }$. (1)
@@
where $\|\cdot\|_{F}$ denotes the Frobenius norm. These problems are both convex, since they minimize a convex function over a convex set. While ($\mathcal{P}$) is well-known in the literature and has been studied in many works [5, 23, 15, 8, 13], ($\mathcal{NP}$) is more recent and has been considered in [15, 14, 4]. Both PSDP and NSPSDP problems arise in the estimate of local compliance matrices during deformable object in various engineering applications; see [14, 4]. The compliance matrix represents the Green function matrix that corresponds to the unknown $A$ in ($\mathcal{NP}$). The estimation of $A$ is based on the observations of the forces, stored in the columns of $X$, and their associated displacements, stored in the columns of $B$, for a total of $m$ measurements. When the deformable object is a passive object, that is, it does not generate energy in deformation, the physical properties of the model impose that $p^{\top}Ap\geq 0$, where $p$ is the point load, which is implied by $A\in\mathcal{N}^{n}_{\succeq}$. Thus, when estimating $A$ under noisy measurements, it is crucial to impose the constraint $A\in\mathcal{N}^{n}_{\succeq}$ in the optimization problem; see [16] and [14, Section III] for more details about the physical model.
@@ 1.1 Contributions and outline of the paper
For both ($\mathcal{P}$) and ($\mathcal{NP}$), the original $n\times m$ dimensional problem can be reduced to a problem of dimension $r\times r$, where $r$ is the rank of $X$; see [8] for ($\mathcal{P}$), and [4] for ($\mathcal{NP}$). This is referred to as a semi-analytical approach because the smaller $r\times r$ dimensional problem still needs to be solved with iterative methods, and [8, 4] relied on an optimal first-order method with linear convergence. It is important to note that the solution of ($\mathcal{P}$) is not always attained [8]; see Section 2.1 where we recall this result. The approach in [4] follows closely that of [8], and concludes that the solution of ($\mathcal{NP}$) is also not always attained. However, it turns out this result is not correct, unless one uses an additional constraint within the NSPSDP problem, which is not explicitly stated in [4].
@@
The paper is organized as follows. In Section 2, we recall the semi-analytical methods proposed in [8] for solving ($\mathcal{P}$) and in [4] for solving ($\mathcal{NP}$). We discuss in detail the main issue of the latter approach, namely the introduction of an unnecessary additional constraint in the problem, which leads to non-optimal and unbounded solutions. In Section 3, we show how to fix this issue, and prove that the infimum of the NSPSDP problem is always attained. We also discuss how to compute the minimum-norm solution and show its low-rank properties. In Section 4, we develop an algorithm relying on the theoretical results from Section 3. Section 5 shows the effectiveness of this new algorithm compared to that of [4] in several numerical examples.
@@ 2.1 The PSDP problem
Before doing this, let us recall an important result about positive semidefinite characterization that will be used throughout the paper.
@@
[23, Theorem 2.2] Let $X,B\in\mathbb{R}^{n\times m}$ and assume that $X$ has rank $n$. Then the infimum of problem ($\mathcal{P}$) is attained for a unique solution $A_{\star}\in\mathcal{S}_{\succeq}^{n}$.
@@
    Otherwise the infimum is not attained, and we can construct a solution $A_{\star}^{(\varepsilon)}\in\mathcal{S}_{\succeq}^{n}$ with minimal rank, minimum Frobenius norm and minimum spectral norm, such that
@@
    for any $\varepsilon>0$ sufficiently small.
@@
We refer to [8, Theorem 1] and [8, Corollary 1] for the proof. ∎
Remark 1.
The explicit expressions for $\varepsilon$ and $A_{\star}^{(\varepsilon)}$ in Theorem 2.3 in the case the infimum is not attained can be found in [8]. These formulas are analogous to the ones of Theorem 2.5 presented in the next section.
@@ 2.2 The NSPSDP problem
The result of Theorem 2.3 for problem ($\mathcal{P}$) has been adapted by Baghel et al. [4] for ($\mathcal{NP}$), but the approach they proposed contains some incorrect statements. Before showing the issues in the main result of [4], we start by recalling an analogous result to Lemma 2.2 for the NSPSDP problem, which guarantees a unique solution when $X$ has full row rank.
@@
[15, Theorem 2.4.6] Let $X,B\in\mathbb{R}^{n\times m}$ and assume that $X$ has rank $n$. Then the infimum of problem ($\mathcal{NP}$) is attained for a unique solution $A_{\star}\in\mathcal{N}_{\succeq}^{n}$.
@@
which is different than the original NSPSDP problem ($\mathcal{NP}$). Although the infimums of ($\mathcal{NPC}$) and ($\mathcal{NP}$) coincide, the optimal solution of ($\mathcal{NPC}$) might not be attained in some cases, see Theorem 2.5 below, while that of ($\mathcal{NP}$) always is; see Theorem 3.1 in the next section.
@@
    Since the infimum of ($\mathcal{NP}$) and ($\mathcal{NPC}$) coincide (see (4)), $A_{\star}$ is also an optimal solution of ($\mathcal{NP}$).
@@
    Otherwise the infimum of ($\mathcal{NPC}$) is not attained, and we can construct a solution $A_{\star}^{(\varepsilon)}\in\mathcal{N}_{\succeq}^{n}$ such that
    $\|A_{\star}^{(\varepsilon)}X-B\|^{2}_{F}\leq\inf_{A\in\mathcal{N}_{\succeq}^{n}}\|AX-B\|^{2}_{F}+\varepsilon$
@@
${\color[rgb]{0,0,0}A_{\star}=U\left(\left(\begin{matrix}\hat{H}_{11}&\frac{1}{2}Z^{\top}\\ \frac{1}{2}Z&K\\ \end{matrix}\right)+\left(\begin{matrix}\hat{S}_{11}&-\frac{1}{2}Z^{\top}\\ \frac{1}{2}Z&R\\ \end{matrix}\right)\right)U^{\top}=U\left(\begin{matrix}\hat{A}_{11}&0\\ Z&K+R\\ \end{matrix}\right)U^{\top},}$
@@
Besides solving a different NSPSDP problem with the additional constraint $A\in\mathcal{N}_{\succeq}^{U\mathcal{C}U^{\top}}$, the issue of the method by Baghel et al. concerns the norm of the solution found when the infimum of ($\mathcal{NPC}$) is not attained, that is, when $X$ has low-rank and $\ker(\hat{H}_{11})\neq\{0\}$ is not contained in $\ker(Z)$. In that case, the matrix $K^{(\varepsilon)}\succeq\frac{1}{4}Z(\hat{H}^{(\varepsilon)}_{11})^{-1}Z^{\top}$ used to define $A_{\star}^{(\varepsilon)}$ has a very large norm, going to infinity as $\varepsilon$ goes to zero. This leads to a huge norm solution, as it will be shown in the numerical examples.
3 A new approach for the NSPSDP problem
@@
the infimum of the NSPSDP problem ($\mathcal{NP}$) is always attained.
We characterize the set of minimizers of the problem and we determine the family of the solutions for the original NSPSDP problem.
@@
$\ker(\hat{H}_{11})\subseteq\ker(Z+M),\qquad N\succeq\frac{1}{4}(Z+M)H_{11}^{\dagger}(Z+M)^{\top},$
so that $A_{\star}\in\mathcal{N}_{\succeq}^{n}$.
@@
Now that we have introduced a new family of solutions for the NSPSDP problem ($\mathcal{NP}$), we provide a simplified characterization of the matrix family $\mathcal{F}$, in which the constraint concerning the kernels in (9) is removed. In this way it will be also easier to highlight low-norm and low-rank properties of matrices in the family.
@@
With the same notation as in Theorem 3.1, the set $\mathcal{F}$ that can be used to characterize optimal solutions of ($\mathcal{NP}$) (see (10)) is given by
@@ 3.1 Minimum-norm solution for the NSPSDP problem
$A=U\begin{pmatrix}\hat{A}_{11}&(YW_{1}^{\top}-Z)^{\top}\\ Z&N+G\\ \end{pmatrix}U^{\top}\in U\mathcal{F}U^{\top}$
where all the matrices fulfil the conditions in Lemma 3.2. Let us introduce the differentiable function
$f(Y):=\underbrace{\|YW_{1}^{\top}-Z\|_{F}^{2}}_{=:g(Y)}+\frac{1}{16}\underbrace{\left\|Y\Lambda^{-1}Y^{\top}\right\|_{F}^{2}}_{=:h(Y)},\qquad Y\in\mathbb{R}^{(n-r)\times s}.$ (11)
Note that $f$ implicitly depends on the fixed matrices $X$ and $B$, since $W_{1}$, $Z$ and $\Lambda$ do.
Lemma 3.3.
The function $f$ defined in (11) is strictly convex in $Y$ and its minimum is unique.
Proof.
We first show that $g$ is strictly convex and that $h$ is convex, therefore their linear combination $f$ is strictly convex. Then we prove that $f$ attains a minimum, which is unique by strict convexity.
Let us define $g_{1}(x)=x^{2}$ and $g_{2}(Y)=\|YW_{1}^{\top}-Z\|_{F}$. The function $g_{1}$ is strictly convex and increasing on $[0,+\infty)$, while the function $g_{2}$ is convex on $\mathbb{R}^{(n-r)\times s}$, since for any $\alpha\in(0,1)$ and any $Y_{1},Y_{2}\in\mathbb{R}^{(n-r)\times s}$,
$g_{2}(\alpha Y_{1}+(1-\alpha)Y_{2})=\|\alpha(Y_{1}W_{1}^{\top}-Z)+(1-\alpha)(Y_{2}W_{1}^{\top}-Z)\|_{F}\leq\alpha g_{2}(Y_{1})+(1-\alpha)g_{2}(Y_{2}).$
Thus $g=g_{1}\circ g_{2}$ is strictly convex. In fact, its Hessian matrix is the identity, meaning that it is perfectly conditioned. The unique minimizer of $g(Y)$ is given by $ZW_{1}$.
Let us now prove the convexity of the quartic function $h(Y)=\|Y\Lambda^{-1}Y^{\top}\|_{F}^{2}=\textnormal{trace}((Y\Lambda^{-1}Y^{\top})^{2})$. Define the diagonal matrix $D:=\Lambda^{-1}$, which has, by construction, positive diagonal entries, and let
$h_{1}(X)=\textnormal{trace}(X^{2}),\qquad h_{2}(Y)=YDY^{\top}.$
The function $h_{1}$ is operator monotone (see [7, Definition 2.1]), that is, for any $X_{1},X_{2}\in\mathcal{S}_{\succeq}^{n-r}$ such that $X_{1}\succ X_{2}$, we have $h_{1}(X_{1})>h_{1}(X_{2})$. Indeed, if $\lambda_{i}$ and $\mu_{i}$ are the eigenvalues of $X_{1}$ and $X_{2}$, respectively,
$\textnormal{trace}(X_{1}^{2}-X_{2}^{2})=\sum_{i=1}^{n-r}\lambda_{i}^{2}-\mu_{i}^{2}=\sum_{i=1}^{n-r}(\lambda_{i}+\mu_{i})(\lambda_{i}-\mu_{i})\geq(n-r)\min_{i}(\lambda_{i}+\mu_{i})\textnormal{trace}(X_{1}-X_{2})>0$
since $X_{1}-X_{2}\succ 0$. Moreover $h_{1}$ is a convex function, since it fulfills the hypothesis of [7, Theorem 2.10], which states that if $\varphi:\mathbb{R}\rightarrow\mathbb{R}$ is convex, then $X\rightarrow\textnormal{trace}(\varphi(X))$ is convex, where $\varphi(X)$ is meant as a matrix function; in this case $\varphi(x)=x^{2}$. The function $h_{2}$ is operator convex, that is, for any $\alpha\in(0,1)$ and any $Y_{1},Y_{2}\in\mathbb{R}^{(n-r)\times s}$,
$\alpha g_{2}(Y_{1})+(1-\alpha)g_{2}(Y_{2})-g_{2}(\alpha Y_{1}-(1-\alpha)Y_{2})$
$=\alpha(1-\alpha)\left(Y_{1}DY_{1}^{\top}+Y_{2}DY_{2}^{\top}-Y_{1}DY_{2}^{\top}-Y_{2}DY_{1}^{\top}\right)=\alpha(1-\alpha)(Y_{1}-Y_{2})D(Y_{1}-Y_{2})^{\top}\succeq 0.$
Thus, for any $\alpha\in(0,1)$ and any $Y_{1},Y_{2}\in\mathbb{R}^{(n-r)\times s}$, the properties of $h_{1}$ and $h_{2}$ yield the convexity of $h$:
$h(\alpha Y_{1}+(1-\alpha)Y_{2})=h_{1}(h_{2}(\alpha Y_{1}+(1-\alpha)Y_{2}))\leq h_{1}(\alpha h_{2}(Y_{1})+(1-\alpha)h_{2}(Y_{2}))\leq\alpha h(Y_{1})+(1-\alpha)h(Y_{2}).$
The function $f$ is trivially bounded from below by 00, and it is continuous with bounded level sets, since for any $\rho>0$
$f(Y)\leq\rho\quad\Rightarrow\quad\|Y\|_{F}=\|YW_{1}^{\top}\|_{F}\leq\|YW_{1}^{\top}-Z\|_{F}+\|Z\|_{F}\leq\sqrt{\rho}+\|Z\|_{F}.$
Thus its infimum is attained while strict convexity implies its uniqueness. ∎
Thanks to the properties of $f(Y)$ shown in Lemma 3.3, the following result provides the unique minimum Frobenius norm solution of ($\mathcal{NP}$), and it also shows its low-rank properties.
@@
Using the same notation as in Theorem 3.1, Lemma 3.2 and Lemma 3.3, the minimum Frobenius norm solution of ($\mathcal{NP}$) is
$A_{\star}:=U\begin{pmatrix}\hat{A}_{11}&(Y_{\star}W_{1}^{\top}-Z)^{\top}\\ Z&\frac{1}{4}Y_{\star}\Lambda^{-1}Y_{\star}^{\top}\\ \end{pmatrix}U^{\top},$
where $Y_{\star}$ is the unique global minimizer of
$\min_{Y\in\mathbb{R}^{(n-r)\times s}}\|YW_{1}^{\top}-Z\|_{F}^{2}+\frac{1}{16}\|Y\Lambda^{-1}Y^{\top}\|_{F}^{2}=\min_{Y\in\mathbb{R}^{(n-r)\times s}}f(Y).$
@@
Since the Frobenius norm is unitary invariant and by using the characterization of Lemma 3.2, we obtain
@@
Hence, minimizing the norm of $A\in U\mathcal{F}U^{\top}$ is equivalent, up to constant summands, to minimize $f(Y)$. Lemma 3.3 shows that the minimum of $f$ is uniquely attained and so $A_{\star}$ is the unique minimizer of $\min_{A\in U\mathcal{F}U^{\top}}\|A\|_{F}^{2}$.
The symmetric part of $A_{\star}$ is